The level of activity at the volcano has been at about the same as yesterday with some rockfalls beginning to produce ash clouds. The number of long-period events have decreased but there has been an increase in volcano-tectonic earthquakes. Good views were obtained of the crater area for most of the day.
Twenty-one volcano-tectonic earthquakes were recorded during the period, an increase over the previous period. These events were all located under the crater at depths of less than 5 km with a few within the upper levels of the volcano. Seven rockfall signals were also recorded; the largest at 11:16 on 18 October, produced a small ash cloud which was observed from the Observatory. One long-period earthquake which was followed by a rockfall was also recorded. Intermittent broadband tremor was recorded on the Gages seismic station from 03:27 to 05:12 today.
Very good visibility for most of the day allowed clear views of the crater area. The new dome has grown and is now just higher than the northern rim of the scar. Expansion of the dome on the floor of the scar produced by the 17/18 explosion has continued towards the west. The eastern face of the dome is now quite steep with a large overhanging block near the rim which looks quite unstable. The back west-north-west side of the dome is now slightly higher than the central areas. Small near-continuous rockfalls were observed from an area on the eastern side of the dome. Small rockfalls were also observed on the western and northern side of the dome.
No Sulphur dioxide measurements was possible today due to unfavourable weather conditions. Processing of the results from yesterday's measurements indicate an sulphur dioxide flux of 1110 tons per day. These results are higher than previous readings and would be closely monitored during next few days.
The sulphur dioxide flux measured by the COSPEC is indicative of sulphur dioxide emission in the volcanic plume and not of gas levels in residential areas. The levels in residential areas are measured by SO2 tubes which have been distributed in occupied areas around the volcano. The gas concentration obtained from these tubes continue to indicate that there is no reason for concern in terms of public health.
EDM measurements were conducted on the eastern triangle to Castle Peak today. The results from this occupation would be available tomorrow.
A survey of the dome was conducted today using the GPS and range finding binoculars. The results of this survey would be compared with the previous survey conducted on 09 October so as to obtain a measure of the magma effusion rates.
The eastern flank of the new dome appears very unstable and it is expected that larger rockfalls and possibly pyroclastic flows may occur during the next few days. The volcano remains in an active and highly dangerous state. Further activity, possibly leading to an explosive eruption, could occur with little warning. Everyone who enters the evacuated zone must remain alert and be ready to move at short notice. Individuals who go beyond the Long Ground area into the Tar River valley are risking death.